40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 15 10:52:38 2019, gautam, Update, SUS, EY door opened 
    Reply  Tue Jan 15 15:49:47 2019, gautam, Update, SUS, EY door opened EY_OSEMs.png
       Reply  Sat Jan 19 15:33:18 2019, gautam, Update, SUS, ETMY OSEMs faulty Screen_Shot_2019-01-19_at_3.32.35_PM.pngETMY_sensors_1231832635.pdf
          Reply  Tue Jan 22 20:36:53 2019, gautam, Update, SUS, ETMY OSEMs faulty 
             Reply  Wed Jan 23 12:39:18 2019, gautam, Update, SUS, EY chamber work 6x
                Reply  Wed Jan 23 23:12:44 2019, gautam, Update, SUS, Prep for FC cleaning IY_level_before.pdf
                   Reply  Thu Jan 24 15:32:31 2019, gautam, Update, SUS, Y arm cavity side first contact applied Yarm_FC.pdfOSEMfilter.png
                      Reply  Tue Jan 29 22:12:40 2019, gautam, Update, SUS, Alignment prep 
                         Reply  Wed Jan 30 11:54:24 2019, gautam, Update, SUS, More alignment prep IMG_5958.JPGIMG_5962.JPG
                            Reply  Wed Jan 30 19:25:40 2019, gautam, Update, SUS, Xarm cavity alignment Xarm.pngXarm_IR.png
                               Reply  Fri Feb 1 01:24:06 2019, gautam, Update, SUS, Almost ready for pumpdown tomorrow 
                                  Reply  Fri Feb 1 13:16:50 2019, gautam, Update, SUS, Pumpdown 83 underway 
                                     Reply  Fri Feb 1 21:52:57 2019, gautam, Update, SUS, Pumpdown 83 underway 
                                  Reply  Fri Feb 1 14:44:14 2019, gautam, Update, SUS, Y arm FC cleaning and reinstall IMG_5974.JPGIMG_5986.JPGIMG_5992.JPG
    Reply  Wed Jan 16 16:25:25 2019, gautam, Update, SUS, Yarm locked Yarm_locked.png
       Reply  Fri Jan 18 21:34:18 2019, gautam, Update, SUS, Unused optic on EY table IMG_5408.JPG
          Reply  Sat Jan 19 05:07:45 2019, Koji, Update, SUS, Unused optic on EY table 
Message ID: 14403     Entry time: Wed Jan 16 16:25:25 2019     In reply to: 14399     Reply to this: 14407
Author: gautam 
Type: Update 
Category: SUS 
Subject: Yarm locked 

[chub, gautam]

Summary:

Y arm was locked at low power in air.

Details:

  1. ITMY chamber door was removed at ~10am with Bob's help.
  2. ETMY table leveling was found to have drifted significantly (3 ticks on the spirit level, while it was more or less level yesterday, should look up the calib of the spirit level into mrad). Chub moved some weights around on the table, we will check the leveling again tomorrow.
  3. IMC was locked.
  4. TT2 and brass alignemnt tool was used to center beam on ETMY.
  5. TT1 and brass alignment tool was used to center beam on ITMY. We had to do a c1susaux reboot to be able to move ITMY. Usual precautions were taken to avoid ITMX getting stuck.
  6. ETMY was used to make return beam from the ETM overlap with the in-going beam near ITMY, using a holey IR card.
  7. At this point, I was confident we would see IR flashes so I decided to do the fine alignment in the control room.

We are operating with 1/10th the input power we normally have, so we expect the IR transmission of the Y arm to max out at 1 when well aligned. However, it is hovering around 0.05 right now, and the dominant source of instability is the angular motion of ETMY due to the Oplev loop being non-functional. I am hesitant to do in-chamber work without an extra pair of eyes/hands around, so I'll defer that for tomorrow morning when Chub gets in. With the cavity axis well defined, I plan to align the green beam to this axis, and use the two to confirm that we are well clear of the Parabola.

* Paola, our vertex laptop, and indeed, most of the laptops inside the VEA, are not ideal to work on this kind of alignmment procedure, it would be good to set up some workstations on which we can easily interact with multiple MEDM screens,

Attachment 1: Yarm_locked.png  74 kB  | Hide | Hide all
Yarm_locked.png
ELOG V3.1.3-